# Passive Beam Shaping via Binary-Coded Apertures

### v2

@@ Abstract
This paper presents coded-aperture reflector for indoor mmWave coverage enhancement in obstructed or blocked LoS settings.We model the reflecting aperture using an equivalent array-factor formulation, where each passive reflecting cell contributes a reradiated field with phase set by the incident and departure directions. Building on this model, we develop two fabrication-friendly passive synthesis methods: (i) binary (1-bit) spatial coding that enables deterministic non-specular beam formation and multi-beam patterns by selecting cell participation on a dense $\lambda/2$ lattice via an ON/OFF metallization mask, and (ii) diffraction-order (periodic) steering that exploits aperture periodicity to place selected diffraction orders at prescribed angles. We analytically characterize the proposed cosine-threshold quantization rule, including its asymptotic activation ratio and a distribution-free lower bound on non-specular gain relative to ideal continuous-phase control. To validate the proposed designs, we fabricate and metallize low-cost prototypes in-house using a copper-backed 3D-printed “inkwell” substrate with stencil-guided conductive-ink deposition. $60$ GHz over-the-air measurements show non-specular power enhancements on the order of +14+142020 dB relative to passive, non-engineered (all-ON) reflector baselines. Results also demonstrate that fully passive, binary-coded apertures can deliver beam control with rapid in-lab manufacturability and offer a practical alternative to power-consuming reconfigurable surfaces for static indoor mmWave links.
@@ I Introduction
A prominent environment-shaping approach is the reconfigurable intelligent surface (RIS), which employs tunable unit cells (e.g., varactors, PIN diodes, MEMS, or liquid crystals) to control the reflection coefficient across an aperture and redirect incident wavefronts [3, 7, 6, 8, 9]. RIS can provide adaptivity, multi-user beam management, and programmable wavefront control, and has therefore received substantial attention in the communications and antennas communities. However, practical RIS deployments must contend with several non-idealities and overheads that are often overlooked in idealized phase-only models. First, finite-resolution phase tuning (quantization), mutual coupling, and element loss can reduce the realized array gain and distort sidelobe structure, especially at mmWave where fabrication tolerances and parasitics matter [9, 8]. Second, RIS operation typically relies on a control loop (CSI acquisition, optimization, and configuration), which can impose signaling overhead and latency, and can become challenging when the channel changes or when users are mobile [10]. Third, and central to the deployment perspective of this work, practical RIS require non-negligible DC power for controllers, driver/bias circuitry, and tunable elements, even if the RF scattering mechanism itself does not include a transmit chain [7, 12, 11]. Measurement-backed studies show that controller/driver power can be at the watt level and that unit-cell dissipation can be state dependent (e.g., for PIN-diode RIS, ON-state cells consume additional power) [12, 11]. For static indoor installations, these operational burdens translate into wiring complexity, power provisioning, maintenance considerations, and cost factors that matter in real deployments but are not captured by the usual “nearly-passive” narrative.
@@
This work focuses on binary-coded apertures that generate deterministic non-specular beams using extremely simple, and strictly passive patterns on a fixed dense lattice. We model the reflecting aperture using an equivalent array-factor formulation, in which each passive reflecting cell contributes a reradiated field whose phase is determined by the incident and departure directions as shown in Fig. 1. Building on this model, we develop two fabrication-friendly passive synthesis methods. First, we propose a closed-form cosine-threshold binary rule that converts the ideal continuous phase ramp for one or more target directions into a fabrication-ready ON/OFF metallization mask on a dense $\lambda/2$ lattice. This binary (1-bit) spatial coding enables beam steering and multi-beam formation by selecting which cells participate in coherent addition. Second, we present diffraction-order (periodic) steering, which exploits controlled aperture periodicity (effective spacing) to place selected diffraction orders at prescribed angles. Both methods are strictly passive and require no bias network, no tuning circuitry, and no surface-side power.
@@
To validate the proposed theory-to-hardware pipeline, we fabricate prototypes using a copper-backed 3D-printed “inkwell” substrate with stencil-guided conductive-ink deposition. Beam patterns are realized by swapping stencils (binary masks) on a base, enabling rapid iteration without redesigning per-cell height profiles. Over-the-air (OTA) measurements at $60$ GHz in single- and multi-beam configurations closely match the predicted steering behavior, confirming that fully passive, binary-coded apertures can deliver deterministic non-specular beam formation with essentially zero operational power and deployment-friendly manufacturability.
@@
  • We validate single- and multi-beam operation through $60$ GHz over-the-air measurements, showing enhancements at intended non-specular directions consistent with theory for both mask-coded and diffraction-order designs.
@@ II-A System Model
We consider a narrowband mmWave link in which a transmitting node illuminates a planar, fully passive reflecting aperture that reradiates energy into the reflection half-space. For analytical tractability, we adopt a one-dimensional (1-D) aperture model along the $x$-axis with $M$ electrically reflecting cells (elements) located at
@@
where $d_{0}$ is the nominal lattice spacing. A dominant plane-wave component impinges on the reflector from the azimuth angle of arrival (AoA) $\theta_{I}$ (measured from aperture broadside), and the scattered field is observed toward an azimuth angle of departure (AoD) $\theta$.
Under the standard far-field array-factor abstraction commonly used for RIS/reflectarrays [6, 7, 26], the complex baseband response can be written in matrix form as
@@
$k=2\pi/\lambda$, and $\mathbf{\Psi}=\mathrm{diag}(\psi_{0},\ldots,\psi_{M-1})$ captures the element-wise complex scattering coefficients. Expanding (2) yields the phasor summation
@@ II-B Problem Formulation
This paper targets blocked-/obstructed-LoS indoor geometries in which extending coverage requires an engineered non-specular reflected path. Given AoA $\theta_{I}$, the design objective is to choose a strictly passive aperture configuration that increases the response toward one or more desired departure angles $\{\theta_{T,\ell}\}_{\ell=1}^{L}$. Equivalently, we seek passive design variables embedded in $\{\psi_{m}\}$ that shape the angular power pattern $|p(\theta,\theta_{I})|^{2}$ to concentrate energy at the target direction(s). Using the scattering response in (4), we seek passive design parameters that shape the angular power pattern $|p(\theta,\theta_{I})|^{2}$ such that
@@ II-B1 Fixed-aperture 1-bit ON/OFF spatial masking
where the binary sequence $\{b_{m}\}$ gates the element phasors and reshapes the spatial spectrum of the aperture. This enables non-specular steering and multi-beam synthesis under a strict 1-bit constraint.
@@ II-C Narrowband Assumption and Wideband Considerations
The analytical development in (4) assumes a narrow band model. Over wide bandwidths, the phase progression scales with $k(f)=2\pi/\lambda(f)$, which can cause frequency-dependent beam pointing (beam squint) and target-direction gain variation when the physical reflector geometry (mask/period) is held fixed. This effect is typically more pronounced for diffraction-order designs because their lobe angles satisfy a wavelength-dependent grating condition. We quantify wideband behavior in Section VII by sweeping frequency while keeping the physical design fixed.
@@ III-B2 Cosine-threshold ON/OFF mapping
$b_{m}\triangleq\frac{\hat{w}_{m}+1}{2}\qquad\Longleftrightarrow\qquad b_{m}=\mathbf{1}\{\cos\phi_{m}\geq 0\}.$ (11)
Thus, the mask retains locations whose ideal phases lie in the constructive half-plane and suppresses those that would contribute destructively at $\theta_{T}$.
@@ III-C Thinning Ratio (ON Fraction)
Lemma 1 (Asymptotic $50\%$ thinning ratio for the cosine-threshold mask)
@@ III-D Distribution-Free Target-Lobe Gain Bound
Lemma 2 (Distribution-free gain bound for ON/OFF masking)
@@
Fig. 3(b) illustrates Lemma 2 by comparing the normalized target-lobe gain of ON/OFF masking with bipolar ($\pm 1$) and ideal continuous-phase control. As expected, $\pm 1$ control incurs a smaller loss (about $3.9$ dB) because sign reversals re-phase contributions that ON/OFF masking must suppress; a similar $3.9$ dB 1-bit quantization loss was reported in [29]. In contrast to nearest-neighbor phase quantization, the present work yields a closed-form 1-bit assignment in which the $\pm 1$ coefficients follow the cosine-sign rule. The figure also confirms the worst-case ON/OFF loss of $9.94$ dB established by Lemma 2. Finally, at $\sin\theta_{T}+\sin\theta_{I}=0$, all schemes achieve $\gamma=1$ (0 dB), corresponding to the specular-reflection condition.
@@ IV Diffraction-Order (Grating-Lobe) Beam Steering via Uniform Period Selection
Section III introduced fixed-aperture ON/OFF masking on a dense $d_{0}\approx\lambda/2$ scaffold. Here we present a complementary passive mechanism based on uniform-period activation. The key design parameter is the inter-element spacing (period) $\delta$ between adjacent active reflecting elements, with $\delta\geq d_{0}$. In practice, $\delta$ can be realized by activating elements with a periodic pattern on the dense lattice (e.g., activating every $p$th lattice location along the synthesis dimension so that $\delta=\kappa d_{0}$). A uniformly periodic aperture produces a discrete set of reflected lobes indexed by an integer diffraction order $n$. By selecting $(\delta,n)$, a chosen order (grating lobe) can be placed at a desired target departure angle using closed-form relationships.
@@ IV-B1 Closed-form period for steering to a target AoD
To align order $n$ with a desired AoD $\theta_{T}$, enforce $\theta_{n}=\theta_{T}$ in (24), yielding
@@ IV-C Multi-Beam Operation via Multiple Visible Diffraction Orders
This section describes the low-cost fabrication workflow used to realize and experimentally validate the two fully passive beam-shaping mechanisms developed in this paper, namely (i) fixed-aperture 1-bit spatial masking (Section III) and (ii) diffraction-order (uniform-period) steering (Section IV). Bot designs share a common 3D-printed “inkwell” substrate that discretizes a $90\times 90~\mathrm{mm}$ aperture into a $35\times 35$ candidate grid ($1225$ wells) with pitch $d_{0}=2.5$ mm (approximately $\lambda/2$ at $60$ GHz). While the synthesis in Section III is developed using a 1-D array-factor model (azimuthal steering), the prototypes are realized as a 2-D lattice to increase effective aperture area and to spatially average fabrication nonidealities (e.g., conductive-ink thickness variation and partial well filling). The common CAD/STL dimensions are summarized in Table I.

### v3

@@ Abstract
This paper presents coded-aperture reflector for indoor millimeter-wave (mmWave) coverage enhancement in obstructed or blocked line-of-sight (LoS) settings.We model the reflecting aperture using an equivalent array-factor formulation, where each passive reflecting cell contributes a reradiated field with phase set by the incident and departure directions. Building on this model, we develop two fabrication-friendly passive synthesis methods: (i) binary (1-bit) spatial coding that enables deterministic non-specular beam formation and multi-beam patterns by selecting cell participation on a dense lattice via an ON/OFF metallization mask, and (ii) diffraction-order (periodic) steering that exploits aperture periodicity to place selected diffraction orders at prescribed angles. We analytically characterize the proposed cosine-threshold quantization rule, including its asymptotic activation ratio and a distribution-free lower bound on non-specular gain relative to ideal continuous-phase control. To validate the proposed designs, we fabricate and metallize low-cost prototypes in-house using a copper-backed 3D-printed “inkwell” substrate with stencil-guided conductive-ink deposition. $60$ GHz over-the-air measurements show non-specular gain enhancements on the order of +14+142020 dB relative to passive, non-engineered (all-ON) reflector baselines. Results also demonstrate that fully passive, binary-coded apertures can deliver beam control with rapid in-lab manufacturability and offer a practical alternative to power-consuming reconfigurable surfaces for static indoor mmWave links.
@@ I Introduction
A prominent environment-shaping approach is the reconfigurable intelligent surface or large intellgent surface (RIS or LIS), which employs tunable unit cells (e.g., varactors, positive-intrinsic-negative (PIN) diodes, microelectromechanical systems (MEMS), or liquid crystals) to control the reflection coefficient across an aperture and redirect incident wavefronts [3, 7, 6, 8, 9]. RIS can provide adaptivity, multi-user beam management, and programmable wavefront control, and has therefore received substantial attention in the communications and antennas communities. However, practical RIS deployments must contend with several non-idealities and overheads that are often overlooked in idealized phase-only models. First, finite-resolution phase tuning (quantization), mutual coupling, and element loss can reduce the realized array gain and distort sidelobe structure, especially at mmWave where fabrication tolerances and parasitics matter [9, 8]. Second, RIS operation typically relies on a control loop (CSI acquisition, optimization, and configuration), which can impose signaling overhead and latency, and can become challenging when the channel changes or when users are mobile [10]. Third, and central to the deployment perspective of this work, practical RIS require non-negligible DC power for controllers, driver/bias circuitry, and tunable elements, even if the RF scattering mechanism itself does not include a transmit chain [7, 12, 11]. Measurement-backed studies show that controller/driver power can be at the watt level and that unit-cell dissipation can be state dependent (e.g., for PIN-diode RIS, ON-state cells consume additional power) [12, 11]. For static indoor installations, these operational burdens translate into wiring complexity, power provisioning, maintenance considerations, and cost factors that matter in real deployments but are not captured by the usual “nearly-passive” narrative.
@@
A practical gap remains between simple passive reflectors, which are robust and power-free but largely limited to specular redirection, and reconfigurable surfaces, which provide flexible beam control at the cost of active hardware, control circuitry, and operational power. Geometry-coded passive reflectors offer another alternative, but they typically require a new 3-D design for each beam objective. To address this gap, this paper proposes a novel copper-backed 3D-printed inkwell reflector platform that combines a fixed dense scaffold with analytically derived binary metallization patterns. This structure enables rapid realization of different beam patterns on a common base reflector design, without requiring a new 3-D reflector geometry for each target response.
Within this framework, this work focuses on binary-coded apertures that generate deterministic non-specular beams using simple passive coding patterns. The reflecting aperture is modeled using an equivalent array-factor formulation, in which each passive reflecting cell contributes a reradiated field whose phase is determined by the incident and departure directions, as shown in Fig. 1. Building on this model, two fabrication-friendly passive synthesis methods are developed. First, a closed-form cosine-threshold rule is proposed to convert the ideal continuous phase ramp for one or more target directions into a fabrication-ready ON/OFF metallization mask on a dense $\lambda/2$ lattice. This binary (1-bit) spatial coding enables beam steering and multi-beam formation by selecting which cells participate in coherent addition. Second, diffraction-order (periodic) steering is developed by exploiting controlled aperture periodicity to place selected diffraction orders at prescribed angles. Both methods are strictly passive and require no bias network, tuning circuitry, or surface-side power.
@@
To validate the proposed theory-to-hardware pipeline, prototypes are fabricated using a copper-backed 3D-printed “inkwell” substrate with stencil-guided conductive-ink deposition. The reflector realization combines a fixed inkwell scaffold, an analytically derived binary ON/OFF mask pattern, and a stencil used to deposit conductive ink only into the selected wells. This enables rapid prototyping, since identical inkwell bases can be prepared in advance and metallized according to different mask patterns. Over-the-air (OTA) measurements at 60.48 GHz in single- and multi-beam configurations closely match the predicted steering behavior, confirming that fully passive binary-coded apertures can deliver deterministic non-specular beam formation with essentially zero operational power and deployment-friendly manufacturability.
@@
  • We validate single- and multi-beam operation through $60.48$ GHz over-the-air measurements, showing enhancements at intended non-specular directions consistent with theory for both mask-coded and diffraction-order designs.
@@ II-A System Model
We consider a narrowband mmWave link in which a transmitting node illuminates a planar, fully passive reflecting aperture that reradiates energy into the reflection half-space. The fabricated reflector used in this work is a 2-D lattice of candidate cells. Accordingly, we begin with a general 2-D planar aperture model, and then specialize it to the azimuth-plane case considered in this paper.
Let the reflector consist of $M_{x}\times M_{y}$ electrically reflecting cells located at
$x_{m}=\left(\frac{M_{x}-1}{2}-m\right)d_{x},\qquad m=0,1,\ldots,M_{x}-1,$ (1)
$y_{n}=\left(\frac{M_{y}-1}{2}-n\right)d_{y},\qquad n=0,1,\ldots,M_{y}-1,$ (2)
where $d_{x}$ and $d_{y}$ denote the lattice spacings along the two aperture axes. A dominant plane-wave component impinges on the reflector from direction $(\theta_{I},\phi_{I})$, where $\theta_{I}$ and $\phi_{I}$ denote the polar and azimuth angles, respectively, and the scattered field is observed toward direction $(\theta,\phi)$. Define the directional-cosine components
$u(\theta,\phi)\triangleq\sin\theta\cos\phi,\qquad v(\theta,\phi)\triangleq\sin\theta\sin\phi.$ (3)
Under the standard far-field array-factor abstraction commonly used for RIS/reflectarrays [6, 7, 26], the complex baseband response can be written as
$p(\theta,\phi;\theta_{I},\phi_{I})=\sum_{m=0}^{M_{x}-1}\sum_{n=0}^{M_{y}-1}\psi_{m,n}e^{-jk\left[x_{m}(u+u_{I})+y_{n}(v+v_{I})\right]},$ (4)
where $k=2\pi/\lambda$, $\psi_{m,n}$ denotes the complex scattering coefficient of the $(m,n)$th cell, and
$u=u(\theta,\phi),\quad v=v(\theta,\phi),\quad u_{I}=u(\theta_{I},\phi_{I}),\quad v_{I}=v(\theta_{I},\phi_{I}).$
In this paper, we are primarily concerned with beam shaping in the azimuth plane. Accordingly, the azimuth angles are fixed to zero, i.e., $\phi=\phi_{I}=0$. Under this restriction, (4) reduces to an azimuth-plane cut of the 2-D aperture response, namely
$p(\theta,0;\theta_{I},0)=\sum_{m=0}^{M_{x}-1}\sum_{n=0}^{M_{y}-1}\psi_{m,n}\,e^{-jkx_{m}\left(\sin\theta+\sin\theta_{I}\right)}.$ (5)
Since the phase progression in (5) depends only on the $x$-coordinates, the 2-D response reduces to an equivalent one-dimensional (1-D) azimuthal model. For analytical tractability, we therefore adopt the 1-D formulation along the $x$-axis with $M$ electrically reflecting cells (elements) located at
@@
where $d_{0}$ is the nominal lattice spacing. A dominant plane-wave component impinges on the reflector from the angle of arrival (AoA) $\theta_{I}$ (measured from aperture broadside), and the scattered field is observed toward an angle of departure (AoD) $\theta$.
Under the same far-field array-factor abstraction, and omitting explicit zero-based indexing for clarity, the complex baseband response can be written in matrix form as
@@
and $\mathbf{\Psi}=\mathrm{diag}(\psi_{0},\ldots,\psi_{M-1})$ captures the element-wise complex scattering coefficients. Expanding (7) yields the phasor summation
@@ II-B Problem Formulation
This paper targets blocked-/obstructed-LoS indoor geometries in which extending coverage requires an engineered non-specular reflected path. In the most general 2-D setting, given incident direction $(\theta_{I},\phi_{I})$, the design objective is to choose a strictly passive aperture configuration that increases the response toward one or more desired departure directions $\{(\theta_{T,\ell},\phi_{T,\ell})\}_{\ell=1}^{L}$. Equivalently, we seek passive design variables embedded in $\{\psi_{m,n}\}$ that shape the angular power pattern $|p(\theta,\phi;\theta_{I},\phi_{I})|^{2}$ to concentrate energy at the target direction(s).
In this paper, however, we focus on azimuth-plane beam shaping with $\phi_{I}=\phi=0$. Under this restriction, the design problem reduces to choosing passive design variables embedded in $\{\psi_{m}\}$ that shape the azimuthal power pattern $|p(\theta,\theta_{I})|^{2}$ to concentrate energy at one or more desired departure angles $\{\theta_{T,\ell}\}_{\ell=1}^{L}$. Using the scattering response in (9), we seek passive design parameters that shape the angular power pattern $|p(\theta,\theta_{I})|^{2}$ such that
@@ II-B1 Fixed-aperture 1-bit ON/OFF spatial masking
where the binary sequence $\{b_{m}\}$ gates the element phasors and reshapes the spatial spectrum of the aperture.
@@ II-C Narrowband Assumption and Wideband Considerations
The analytical development in (9) assumes a narrowband model. Over wide bandwidths, the phase progression scales with $k(f)=2\pi/\lambda(f)$, which can cause frequency-dependent beam pointing (beam squint) and target-direction gain variation when the physical reflector geometry (mask/period) is held fixed. Because the physical cell spacing is fixed, its electrical spacing varies with frequency; as frequency increases, the lattice becomes electrically sparser relative to the wavelength, whereas at lower frequencies it becomes electrically denser. This effect is typically more pronounced for diffraction-order designs because their lobe angles satisfy a wavelength-dependent grating condition. We quantify wideband behavior in Section VIII by sweeping frequency while keeping the physical design fixed.
@@ III-B2 Cosine-threshold ON/OFF mapping
$b_{m}\triangleq\frac{\hat{w}_{m}+1}{2}\qquad\Longleftrightarrow\qquad b_{m}=\mathbf{1}\{\cos\phi_{m}\geq 0\},$ (16)
where $\mathbf{1}\{\cdot\}$ denotes the indicator function, i.e., $\mathbf{1}\{\mathcal{A}\}=1$ if the condition $\mathcal{A}$ is true and $\mathbf{1}\{\mathcal{A}\}=0$ otherwise. Thus, the mask retains locations whose ideal phases lie in the constructive half-plane and suppresses those that would contribute destructively at $\theta_{T}$.
@@ III-C Thinning Ratio (ON Fraction)
Lemma 1 (Asymptotic $50\%$ thinning ratio for the single-target cosine-threshold mask)
@@ III-D Distribution-Free Target-Lobe Gain Bound
Lemma 2 (Distribution-free gain bound for single-target ON/OFF masking)
@@
Lemma 2 gives a worst-case guarantee on target-direction performance under the ON/OFF masking constraint. Specifically, the maximum normalized target-lobe gain is always at least $1/\pi^{2}$, which corresponds to a worst-case loss of $9.94$ dB relative to the ideal continuous-phase upper bound. Fig. 3(b) illustrates this result by comparing the normalized target-lobe gain of ON/OFF masking with bipolar ($\pm 1$) and ideal continuous-phase control. As expected, $\pm 1$ control incurs a smaller loss (about $3.9$ dB), since sign reversals can re-phase contributions that ON/OFF masking must suppress; a similar $3.9$ dB 1-bit quantization loss was reported in [29], though this work uses a closed-form 1-bit assignment rule. The figure also confirms the $9.94$ dB worst-case ON/OFF loss established by Lemma 2. Finally, at $\sin\theta_{T}+\sin\theta_{I}=0$, all schemes achieve $\gamma=1$ (0 dB), corresponding to the specular-reflection condition.
@@ IV Diffraction-Order (Grating-Lobe) Beam Steering via Uniform Period Selection
Section III introduced fixed-aperture ON/OFF masking on a dense $d_{0}\approx\lambda/2$ scaffold. Here we present a complementary passive mechanism based on uniform-period activation. The key design parameter is the inter-element spacing (period) $\delta$ between adjacent active reflecting elements, with $\delta\geq d_{0}$. In practice, $\delta$ can be realized by activating elements with a periodic pattern on the dense lattice (e.g., activating every $m$th lattice location along the synthesis dimension so that $\delta=\kappa d_{0}$). A uniformly periodic aperture produces a discrete set of reflected lobes indexed by an integer diffraction order $n$. By selecting $(\delta,n)$, a chosen order (grating lobe) can be placed at a desired target departure angle using closed-form relationships.
@@ IV-B1 Closed-form period for steering to a target AoD
To align order $n$ with a desired AoD $\theta_{T}$, enforce $\theta_{n}=\theta_{T}$ in (29) to obtain
@@ IV-C Multi-Beam Operation via Multiple Visible Diffraction Orders
This section describes the low-cost fabrication workflow used to realize and experimentally validate the two fully passive beam-shaping mechanisms developed in this paper, namely (i) fixed-aperture 1-bit spatial masking (Section III) and (ii) diffraction-order (uniform-period) steering (Section IV). Both designs are implemented on a common 3D-printed “inkwell” reflector platform, in which a planar aperture is discretized into a dense lattice of candidate metallization locations. The reflector realization involves three components. The first is the 3D-printed inkwell scaffold, which serves as the physical base platform. The second is the analytically derived binary mask pattern, which specifies the ON/OFF metallization layout. The third is the stencil, which is used during fabrication to deposit conductive ink only into the selected wells. Since the beam synthesis in this paper is carried out for azimuthal steering with the elevation angle fixed to zero, the fabricated prototypes are realized on a 2-D lattice by replicating the same 1-D azimuthal pattern across rows. The reflector dimensions, lattice parameters, material choices, and operating frequency used in this paper are summarized in Table I.
Fig. 4 shows a representative reflector cell in top and side view. Each cell consists of a conductive ink/metallized fill inside an inkwell cavity, a 3D-printed dielectric scaffold, and a metallic copper backing realized using copper adhesive tape directly attached to the dielectric. This yields a metal-backed dielectric structure similar in structural spirit to prior passive reflector implementations such as MilliMirror [24], although the present work does not rely on thickness-tuned unit-cell phase synthesis.
The fabricated reflectors are designed using the aperture-level analytical array model developed in this paper. In this model, each lattice location is treated as an effective reflector sample contributing to the reradiated field. The active (metallized) locations are selected either by the cosine-threshold rule described in Section III-B or by the periodic-spacing selection described in Section IV-A. These analytical results are used both to design the fabricated reflectors and to generate the corresponding theoretical patterns. Accordingly, this work does not include full-wave extraction of element-level reflection magnitude and phase under normal or oblique incidence, as would be done in a Computer Simulation Technology (CST) or High Frequency Structure Simulator (HFSS)-based metasurface unit-cell analysis. Similar element-/surface-level abstractions are also common in RIS system modeling [6, 7]. In the analytical model, metallized cells are treated as ideal reflecting points, while non-metallized cells are assumed to contribute negligibly to the reflected field. Full-wave element-level reflection characterization under normal and oblique incidence is left for future work.
Figure 4: Representative reflector cell (element) geometry: (a) top view and (b) side view. The cell consists of a conductive ink/metallized fill formed inside the inkwell cavity, a 3D-printed dielectric scaffold, and a metallic copper backing in direct contact with the dielectric.